### v1

@@ Abstract
This work compares three categories of explainability for gate-level hardware trojan detection on the Trust-Hub benchmark dataset: (1) domain-aware property-based analysis of 31 circuit-specific features derived from gate fanin patterns, flip-flop distances, and primary I/O connectivity; (2) model-agnostic case-based reasoning using k-nearest neighbors for precedent-based explanations; and (3) model-agnostic feature attribution methods (LIME, SHAP, gradient) that provide generic importance scores without circuit-level context.
The findings show that different explainability approaches offer distinct advantages for hardware security practitioners. The domain-aware property-based method analyzes 31 circuit properties, enabling engineers to understand detection decisions through familiar concepts like "high fanin complexity near outputs indicates potential trojan triggers." Case-based reasoning (k-nearest neighbors) achieves 97.4% correspondence between predictions and training exemplars, providing model-agnostic justifications grounded in precedent that require minimal domain interpretation. In contrast, while LIME and SHAP provide mathematically rigorous feature attributions with strong inter-method correlation ($\rho$=0.94, p<<0.001), they produce generic feature importance scores that lack circuit-level context for validation or remediation planning. Detection performance using XGBoost classification with optimized threshold achieves 46.15% precision and 52.17% recall on 11,392 test samples, representing a 9-fold improvement over prior work (Hasegawa et al.: 5.13% precision) while reducing false positive rates from approximately 5.6% to 0.25%.
@@
This work provides empirical evidence comparing domain-aware property analysis, model-agnostic case-based reasoning, and model-agnostic feature attribution techniques for hardware security applications. The results show that property-based and case-based approaches offer complementary advantages: domain alignment and precedent-based interpretability, respectively, compared to generic feature rankings. These findings have implications for XAI deployment across domains where practitioners must validate and act on ML predictions.
@@ 1 Introduction
Machine learning approaches have demonstrated promise for hardware trojan detection by identifying suspicious circuit structures through statistical analysis of gate-level netlists. However, existing ML-based detection methods face a challenge: opaque-box models provide predictions without interpretable justifications that hardware security engineers can understand and act upon. While explainable AI (XAI) techniques have been developed to interpret ML predictions, these methods were primarily designed for general domains like image classification and natural language processing. This raises the question: Do model-agnostic explainability techniques developed for general ML applications provide the actionable insights that hardware security practitioners need, or do domain-aware methods tailored to circuit characteristics offer interpretability?
@@
    Domain-Aware Property Analysis: A property-based explainability method is developed and evaluated that analyzes 31 circuit-specific features grounded in hardware design principles. This domain-aware approach provides explanations like "high Logic Gate Fanin (LGFi=12) at 2 levels upstream combined with low Flip-Flop Output distance (ffo=1) matches rare-event trigger patterns," enabling hardware engineers to validate findings through familiar circuit analysis concepts.
@@
    Improved Detection Performance with XGBoost: The XGBoost-based detector achieves 46.15% precision (95% CI: [42.06%, 50.24%]) at 52.17% recall (95% CI: [45.65%, 58.70%]) on Trust-Hub circuits, representing a 9-fold improvement in precision over the Hasegawa et al. baseline (5.13% precision, 82.6% recall). This improvement addresses a key limitation for practical deployment where engineers must investigate flagged gates.
@@ 2.1 Hardware Trojans
Figure 2 shows a specific implementation of a confidentially hardware trojan using logic gates within an IC design file netlist. The dashed lines represent additional levels of logic gates. The trigger is conditional on input patterns within the IC, which eventually originate from the primary inputs. The logical-and gate requires that all the inputs are logically true before triggering the payload. To be difficult to detect or accidentally trigger during IC test mode, the hardware trojan trigger should be a rare event condition. Given $n$ inputs to a combinational IP circuit, the rarest condition would be one of $2^{n}$ inputs. Once the hardware trojan is triggered, the multiplexor circuit within the payload switches from normal operation to leaking sensitive information on the primary output pin of the circuit where the attacker can observe it.
@@
Hasegawa et al. introduced a method of detecting hardware trojans from a gate-level netlist [4, 5]. Five features from each net are used to classify trojans. (a) Logic Gate Fanins (LGFi) represent the number of inputs to the logic gates two levels upstream from a net. (b) Flip-Flop Input (FFi) represents the number of upstream logic levels to a flip-flop. (c) Flip-Flop Output (FFo) represents the number of logic levels downstream to a flip-flop. (d) Primary Input (PI) which represents the number of logic levels upstream to the closest primary input. (e) Primary Output (PO) represents the number of logic levels downstream to the closest primary output. In the work, the authors also used three training strategies, due to the imbalance of the trojan to non-trojan training data ratio: no weighting, static weighting, and dynamic weighting. Dynamic weighting performed best. Other works address the imbalance in other means, such as synthetic method over-sampling techniques [7]. These optimized methods still do not provide a better understanding of the ML results.
2.2 Trust-Hub Trojan Database
The Trust-Hub benchmark repository [8, 9, 10, 6] provides gate-level netlists with known trojan insertions for various circuit types. This work uses 30 circuits comprising RS232 communication peripherals and ISCAS combinational benchmarks.
@@ 3.1 Machine Learning for Hardware Trojan Detection
Hasegawa et al. established a foundational approach for ML-based hardware trojan detection from gate-level netlists [4, 5]. Their method extracts five structural features from each net: Logic Gate Fanins (LGFi), Flip-Flop Input distance (FFi), Flip-Flop Output distance (FFo), Primary Input distance (PI), and Primary Output distance (PO). Using Support Vector Machine (SVM) classification with dynamic class weighting to address data imbalance, they reported 82.6% recall on Trust-Hub circuits.
@@ 3.1.3 Ensemble Methods for Hardware Trojan Detection
Ensemble learning combines multiple classifiers to improve detection performance through model diversity and aggregated predictions. Negishi and Togawa [7] evaluated ensemble learning models for gate-level hardware trojan identification, comparing Random Forest, AdaBoost, and Gradient Boosting on Trust-Hub circuits. Their experimental results demonstrated that ensemble methods can achieve higher F-measures compared to single classifiers like SVM, with Random Forest showing particular effectiveness for imbalanced trojan detection tasks.
@@ 3.1.4 Class Imbalance Handling in Hardware Trojan Detection
Hardware trojan detection faces extreme class imbalance: in typical gate-level netlists, benign gates outnumber trojan gates by ratios exceeding 100:1, often reaching 275:1 as observed in Trust-Hub datasets. Standard machine learning classifiers trained on such imbalanced data exhibit bias toward the majority class, achieving high overall accuracy while failing to detect minority-class trojans.
@@ 3.2.1 General XAI Methods
Local Interpretable Model-Agnostic Explanations (LIME) [1] explains individual predictions by perturbing inputs and fitting interpretable local models. LIME has been applied to image classification and text analysis but has not been systematically evaluated for hardware security applications. Shapley Additive Explanations (SHAP) [2] uses game-theoretic Shapley values to compute feature attributions with theoretical guarantees of consistency and local accuracy. However, SHAP’s computational cost can be prohibitive for large feature spaces or extensive datasets.
@@ 3.2.2 XAI for Hardware Security
3.3 Explainable Artificial Intelligence (XAI)
@@
Case-based explanations for medical models, introduced by Caruana et al., suggested using a method based on k-nearest neighbor (KNN) distance in multidimensional feature space as effective in identifying like cases from training as explanations for new samples [13]. Cases from training should produce results similar to new samples if they are alike. The case-based method suggested that leveraging training data is more difficult for complex models such as NN as the training set is discarded. In the case of NNs, the activation of $n$ hidden neurons is translated into points in an $n$-dimensional space, and a KNN algorithm can be applied to find similar activation patterns. While this may suggest similarity to the NN’s activation and behavior between like cases, the method does little to explain what is happening in the NN.
Outstanding work on Local Interpretable Model-Agnostic Explanations (LIME) explains individual samples by perturbing the sample to retrieve local elements and present them to an explainable model[1]. LIME trains a local linear model by systematically perturbing feature values around the instance being explained and observing how predictions change. For tabular data, LIME perturbs each feature independently by sampling from the training distribution, generating thousands of perturbed samples to approximate the local decision boundary.
@@
Shapley Additive Explanations (SHAP) also provides quantitative explanations for predictions by assigning each feature an importance value based on game-theoretic Shapley values[2]. SHAP computes the contribution of each feature by considering all possible feature coalitions, measuring how much each feature changes the prediction when added to different subsets of other features. The resulting SHAP values sum to the difference between the model’s prediction and the baseline (average) prediction, providing a complete additive explanation.
@@ 4.1 Feature Extraction from Gate-Level Netlists
The five base features are: (1) Logic Gate Fanin (LGFi): number of inputs to logic gates two levels upstream from a gate, capturing input complexity that may indicate trigger circuits requiring rare input combinations (a trigger with $n$ inputs requires one specific combination from $2^{n}$ possibilities); (2) Flip-Flop Input distance (FFi): number of logic levels upstream to the nearest flip-flop, measuring proximity to state elements; (3) Flip-Flop Output distance (FFo): number of logic levels downstream to the nearest flip-flop, indicating signal propagation paths; (4) Primary Input distance (PI): number of logic levels upstream to the nearest primary input; (5) Primary Output distance (PO): number of logic levels downstream to the nearest primary output, relevant for payload propagation to observable locations where attackers can extract sensitive data.
@@ 4.2 XGBoost Classifier for Trojan Detection
XGBoost (eXtreme Gradient Boosting) [28] is used as the binary classifier for gate-level trojan detection. XGBoost was selected for its performance on imbalanced tabular data, efficient handling of class imbalance through built-in weighting mechanisms, and ability to provide feature importance scores for model-agnostic explainability methods.
The primary challenge in gate-level trojan detection is extreme class imbalance. In the Trust-Hub dataset, the ratio of benign gates to trojan gates is approximately 275:1 (11,346 benign gates vs. 46 trojan gates in test data). Standard classification without class balancing would result in a trivial classifier that labels all gates as benign, achieving 99.6% accuracy but 0% recall for trojans.
@@
The XGBoost model uses the following hyperparameters: 100 estimators (decision trees), maximum depth of 5 to prevent overfitting on sparse trojan patterns, learning rate of 0.1, and objective=’binary:logistic’ for binary classification with probability outputs. Training is performed on 80% of the combined Trust-Hub dataset (45,567 samples) with evaluation on the held-out 20% test set (11,392 samples), maintaining the natural class imbalance in both splits to reflect real-world deployment scenarios.
@@ 4.3.2 Inherently Interpretable: Case-Based Reasoning with k-Nearest Neighbors
Case-Based explainability is implemented using the case-explainer package [30], a general-purpose library for model-agnostic case-based explanations with k-Nearest Neighbors (k-NN). The method uses $k=5$ neighbors in the five-dimensional feature space. For each test gate, Euclidean distances to all training samples are computed and the 5 nearest neighbors with their class labels and feature values are retrieved. Predictions are made by majority vote among the $k$ neighbors’ classes.
@@ 5.1 Trust-Hub Benchmark Dataset
Five structural features are extracted from each gate following the methodology of Hasegawa et al. [4, 5]: Local Gate Fanin (LGFi), Flip-Flop Fanin (FFi), Flip-Flop Fanout (FFo), Primary Input distance (PI), and Primary Output distance (PO). Feature extraction is performed by parsing Verilog netlists using CircuitGraph [25] and constructing directed graph representations with NetworkX [27] to query structural properties of each gate.
@@ 5.3.1 Detection Performance Metrics
For binary classification performance, the method report:
@@ 5.3.2 Explainability Evaluation Metrics
Qualitative evaluation through manual inspection of explanation outputs assesses whether explanations provide actionable insights for hardware security engineers.
@@ 5.5.3 Effect Size Reporting
To quantify practical significance beyond statistical significance, the method report standardized effect sizes:
  • Cohen’s d: For continuous metrics, the method compute $d=\frac{\mu_{1}-\mu_{2}}{\sigma_{pooled}}$ where $\sigma_{pooled}$ is the pooled standard deviation. Values of $|d|\geq 0.8$ indicate large effects.
  • Odds ratios: For binary classification improvements, the method report the odds ratio of correct classification between methods, quantifying the multiplicative improvement in odds.
  • Relative improvement: For metrics like false positive rate reduction, the method report both absolute differences (percentage points) and relative improvements (fold reduction).
@@ 6 Results
All experiments were conducted on the Trust-Hub dataset with 56,959 total samples (45,567 training, 11,392 test), derived from 30 circuits containing 46 trojan instances. Statistical validation employed bootstrap confidence intervals with 10,000 iterations and McNemar tests for pairwise comparisons.
@@ 6.2 Statistical Significance
Bootstrap confidence intervals (10,000 iterations, seed=42) confirmed the statistical robustness of the optimized XGBoost method. At the baseline threshold of 0.5, precision confidence interval was [7.14%, 13.53%], demonstrating the instability of unoptimized classification. The optimized threshold (0.99) narrowed this interval to [42.06%, 50.24%], indicating more reliable performance.
McNemar test comparing Method 1 (ensemble) to Method 2 (single XGBoost) found statistically significant differences with $\chi^{2}=1431.25$, p < 0.001. The large chi-square value is partly influenced by the large sample size (n=11,392n=11{,}392). To assess practical significance beyond statistical significance, the effect size is computed: with the observed discordant pairs, the proportions of correct classifications differ between methods (Method 2: 99.56% accuracy vs. Method 1: 84.45%), representing a 15.11 percentage point improvement. Note that Methods 2-5 share identical predictions (same classifier, different explanations), so statistical tests between these methods measure only differences in explainability coverage and computational efficiency, not classification accuracy.

### v2

@@ Abstract
This work compares three categories of explainability for gate-level hardware trojan detection on the Trust-Hub benchmark dataset: (1) domain-aware property-based analysis of 31 circuit-specific features derived from gate fanin patterns, flip-flop distances, and primary I/O connectivity; (2) model-agnostic case-based reasoning using k-nearest neighbors for precedent-based explanations; and (3) model-agnostic feature attribution methods (Local Interpretable Model-agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), gradient) that provide generic importance scores without circuit-level context.
The findings show that different explainability approaches offer distinct advantages for hardware security practitioners. The domain-aware property-based method analyzes 31 circuit properties, enabling engineers to understand detection decisions through familiar concepts like "high fanin complexity near outputs indicates potential trojan triggers." Case-based reasoning (k-nearest neighbors) achieves 97.4% correspondence between predictions and training exemplars, providing model-agnostic justifications grounded in precedent that require minimal domain interpretation. In contrast, while LIME and SHAP provide mathematically rigorous feature attributions with strong inter-method correlation ($\rho$=0.94, p<<0.001), they produce generic feature importance scores that lack circuit-level context for validation or remediation planning. Detection performance using eXtreme Gradient Boosting (XGBoost) classification with optimized threshold achieves 46.15% precision and 52.17% recall on 11,392 test samples, representing a 9-fold improvement over prior work (Hasegawa et al.: 5.13% precision) while reducing false positive rates from approximately 5.6% to 0.25%.
@@
This work provides empirical evidence comparing domain-aware property analysis, model-agnostic case-based reasoning, and model-agnostic feature attribution techniques for hardware security applications. The results show that property-based and case-based approaches offer complementary advantages: domain alignment and precedent-based interpretability, respectively, compared to generic feature rankings. These findings have implications for explainable Artificial Intelligence (XAI) deployment across domains where practitioners must validate and act on Machine Learning (ML) predictions.
@@ 1 Introduction
Machine learning approaches have demonstrated promise for hardware trojan detection by identifying suspicious circuit structures through statistical analysis of gate-level netlists. However, existing ML-based detection methods face a challenge: opaque-box models provide predictions without interpretable justifications that hardware security engineers can understand and act upon. While XAI techniques have been developed to interpret ML predictions, these methods were primarily designed for general domains like image classification and natural language processing. This raises the question: Do model-agnostic explainability techniques developed for general ML applications provide the actionable insights that hardware security practitioners need, or do domain-aware methods tailored to circuit characteristics offer interpretability?
@@
    Domain-Aware Property Analysis: A property-based explainability method is developed and evaluated that analyzes 31 circuit-specific features grounded in hardware design principles. This domain-aware approach provides explanations like "high Logic Gate Fanin (LGFi)=12 at 2 levels upstream combined with low Flip-Flop Output distance (FFo)=1 matches rare-event trigger patterns," enabling hardware engineers to validate findings through familiar circuit analysis concepts.
@@
    Improved Detection Performance with XGBoost: The XGBoost-based detector achieves 46.15% precision (95% CI: [42.06%, 50.24%]) at 52.17% recall (95% CI: [45.65%, 58.70%]) on the Trust-Hub benchmark dataset, representing a 9-fold improvement in precision over the Hasegawa et al. baseline (5.13% precision, 82.6% recall). This improvement addresses a key limitation for practical deployment where engineers must investigate flagged gates.
@@ 2.1 Hardware Trojans
Figure 2 shows a specific implementation of a confidential hardware trojan using logic gates within an Integrated Circuit (IC) design file netlist. The dashed lines represent additional levels of logic gates. The trigger is conditional on input patterns within the IC, which eventually originate from the primary inputs. The logical-and gate requires that all the inputs are logically true before triggering the payload. To be difficult to detect or accidentally trigger during IC test mode, the hardware trojan trigger should be a rare event condition. Given $n$ inputs to a combinational IP circuit, the rarest condition would be one of $2^{n}$ inputs. Once the hardware trojan is triggered, the multiplexor circuit within the payload switches from normal operation to leaking sensitive information on the primary output pin of the circuit where the attacker can observe it.
@@
Hasegawa et al. introduced a method of detecting hardware trojans from a gate-level netlist [4, 5]. Five features from each net are used to classify trojans. (a) Logic Gate Fanins: LGFi represent the number of inputs to the logic gates two levels upstream from a net. (b) Flip-Flop Input (FFi) represents the number of upstream logic levels to a flip-flop. (c) Flip-Flop Output: FFo represents the number of logic levels downstream to a flip-flop. (d) Primary Input (PI) which represents the number of logic levels upstream to the closest primary input. (e) Primary Output (PO) represents the number of logic levels downstream to the closest primary output. In the work, the authors also used three training strategies, due to the imbalance of the trojan to non-trojan training data ratio: no weighting, static weighting, and dynamic weighting. Dynamic weighting performed best. Other works address the imbalance in other means, such as synthetic method over-sampling techniques [7]. These optimized methods still do not provide a better understanding of the ML results.
2.2 Trust-Hub Benchmark Dataset
The Trust-Hub benchmark dataset [8, 9, 10, 6] provides gate-level netlists with known trojan insertions for various circuit types. This work uses 30 circuits comprising RS232 communication peripherals and International Symposium on Circuits and Systems (ISCAS) combinational benchmarks.
@@ 3.1 Machine Learning for Hardware Trojan Detection
Hasegawa et al. established a foundational approach for ML-based hardware trojan detection from gate-level netlists [4, 5]. Their method extracts five structural features from each net: LGFi, FFi, FFo, PI, and PO. Using Support Vector Machine (SVM) classification with dynamic class weighting to address data imbalance, they reported 82.6% recall on the Trust-Hub benchmark dataset.
@@ 3.1.3 Ensemble Methods for Hardware Trojan Detection
Ensemble learning combines multiple classifiers to improve detection performance through model diversity and aggregated predictions. Negishi and Togawa [7] evaluated ensemble learning models for gate-level hardware trojan identification, comparing Random Forest, AdaBoost, and Gradient Boosting on the Trust-Hub benchmark dataset. Their experimental results demonstrated that ensemble methods can achieve higher F-measures compared to single classifiers like SVM, with Random Forest showing particular effectiveness for imbalanced trojan detection tasks.
@@ 3.1.4 Class Imbalance Handling in Hardware Trojan Detection
Hardware trojan detection faces extreme class imbalance: in typical gate-level netlists, benign gates outnumber trojan gates by ratios exceeding 100:1, often reaching 275:1 as observed in the Trust-Hub benchmark dataset. Standard machine learning classifiers trained on such imbalanced data exhibit bias toward the majority class, achieving high overall accuracy while failing to detect minority-class trojans.
@@ 3.2.1 General XAI Methods
LIME [1] explains individual predictions by perturbing inputs and fitting interpretable local models. LIME has been applied to image classification and text analysis but has not been systematically evaluated for hardware security applications. SHAP [2] uses game-theoretic Shapley values to compute feature attributions with theoretical guarantees of consistency and local accuracy. However, SHAP’s computational cost can be prohibitive for large feature spaces or extensive datasets.
@@ 3.2.2 XAI for Hardware Security
3.3 Explainable Artificial Intelligence
@@
Case-based explanations for medical models, introduced by Caruana et al., suggested using a method based on k-nearest neighbor (k-NN) distance in multidimensional feature space as effective in identifying like cases from training as explanations for new samples [13]. Cases from training should produce results similar to new samples if they are alike. The case-based method suggested that leveraging training data is more difficult for complex models such as NN as the training set is discarded. In the case of NNs, the activation of $n$ hidden neurons is translated into points in an $n$-dimensional space, and a k-NN algorithm can be applied to find similar activation patterns. While this may suggest similarity to the NN’s activation and behavior between like cases, the method does little to explain what is happening in the NN.
Outstanding work on LIME explains individual samples by perturbing the sample to retrieve local elements and present them to an explainable model[1]. LIME trains a local linear model by systematically perturbing feature values around the instance being explained and observing how predictions change. For tabular data, LIME perturbs each feature independently by sampling from the training distribution, generating thousands of perturbed samples to approximate the local decision boundary.
@@
SHAP also provides quantitative explanations for predictions by assigning each feature an importance value based on game-theoretic Shapley values[2]. SHAP computes the contribution of each feature by considering all possible feature coalitions, measuring how much each feature changes the prediction when added to different subsets of other features. The resulting SHAP values sum to the difference between the model’s prediction and the baseline (average) prediction, providing a complete additive explanation.
@@ 4.1 Feature Extraction from Gate-Level Netlists
The five base features are: (1) LGFi: number of inputs to logic gates two levels upstream from a gate, capturing input complexity that may indicate trigger circuits requiring rare input combinations (a trigger with $n$ inputs requires one specific combination from $2^{n}$ possibilities); (2) FFi: number of logic levels upstream to the nearest flip-flop, measuring proximity to state elements; (3) FFo: number of logic levels downstream to the nearest flip-flop, indicating signal propagation paths; (4) PI: number of logic levels upstream to the nearest primary input; (5) PO: number of logic levels downstream to the nearest primary output, relevant for payload propagation to observable locations where attackers can extract sensitive data.
@@ 4.2 XGBoost Classifier for Trojan Detection
XGBoost [28] is used as the binary classifier for gate-level trojan detection. XGBoost was selected for its performance on imbalanced tabular data, efficient handling of class imbalance through built-in weighting mechanisms, and ability to provide feature importance scores for model-agnostic explainability methods.
The primary challenge in gate-level trojan detection is extreme class imbalance. In the Trust-Hub benchmark dataset, the ratio of benign gates to trojan gates is approximately 275:1 (11,346 benign gates vs. 46 trojan gates in test data). Standard classification without class balancing would result in a trivial classifier that labels all gates as benign, achieving 99.6% accuracy but 0% recall for trojans.
@@
The XGBoost model uses the following hyperparameters: 100 estimators (decision trees), maximum depth of 5 to prevent overfitting on sparse trojan patterns, learning rate of 0.1, and objective=’binary:logistic’ for binary classification with probability outputs. Training is performed on 80% of the combined Trust-Hub benchmark dataset (45,567 samples) with evaluation on the held-out 20% test set (11,392 samples), maintaining the natural class imbalance in both splits to reflect real-world deployment scenarios.
@@ 4.3.2 Inherently Interpretable: Case-Based Reasoning with k-Nearest Neighbors
Case-Based explainability is implemented using the case-explainer package [30], a general-purpose library for model-agnostic case-based explanations with k-NN. The method uses $k=5$ neighbors in the five-dimensional feature space. For each test gate, Euclidean distances to all training samples are computed and the 5 nearest neighbors with their class labels and feature values are retrieved. Predictions are made by majority vote among the $k$ neighbors’ classes.
@@ 5.1 Trust-Hub Benchmark Dataset
Five structural features are extracted from each gate following the methodology of Hasegawa et al. [4, 5]: LGFi, FFi, FFo, PI, and PO. Feature extraction is performed by parsing Verilog netlists using CircuitGraph [25] and constructing directed graph representations with NetworkX [27] to query structural properties of each gate.
@@ 5.3.1 Detection Performance Metrics
For binary classification performance, the method reports metrics based on True Positives (TP), False Positives (FP), True Negatives (TN), and False Negatives (FN):
@@ 5.3.2 Explainability Evaluation Metrics
Qualitative evaluation through manual inspection of explanation outputs assess whether explanations provide actionable insights for hardware security engineers.
@@ 5.5.3 Effect Size Reporting
To quantify practical significance beyond statistical significance, the method reports standardized effect sizes:
  • Cohen’s d: For continuous metrics, the method computes $d=\frac{\mu_{1}-\mu_{2}}{\sigma_{pooled}}$ where $\sigma_{pooled}$ is the pooled standard deviation. Values of $|d|\geq 0.8$ indicate large effects.
  • Odds ratios: For binary classification improvements, the method reports the odds ratio of correct classification between methods, quantifying the multiplicative improvement in odds.
  • Relative improvement: For metrics like false positive rate reduction, the method reports both absolute differences (percentage points) and relative improvements (fold reduction).
@@ 6 Results
All experiments were conducted on the Trust-Hub benchmark dataset with 56,959 total samples (45,567 training, 11,392 test), derived from 30 circuits containing 46 trojan instances. Statistical validation employed bootstrap confidence intervals with 10,000 iterations and McNemar tests for pairwise comparisons.
@@ 6.2 Statistical Significance
Bootstrap confidence intervals (10,000 iterations, seed=42) confirmed the statistical robustness of the optimized XGBoost method. At the baseline threshold of 0.5, the precision confidence interval was [7.14%, 13.53%], demonstrating the instability of unoptimized classification. The optimized threshold (0.99) narrowed this interval to [42.06%, 50.24%], indicating more reliable performance.
A McNemar test comparing Method 1 (ensemble) to Method 2 (single XGBoost) found statistically significant differences with $\chi^{2}=1431.25$, p < 0.001. The large chi-square value is partly influenced by the large sample size (n=11,392n=11{,}392). To assess practical significance beyond statistical significance, the effect size is computed: with the observed discordant pairs, the proportions of correct classifications differ between methods (Method 2: 99.56% accuracy vs. Method 1: 84.45%), representing a 15.11 percentage point improvement. Note that Methods 2-5 share identical predictions (same classifier, different explanations), so statistical tests between these methods measure only differences in explainability coverage and computational efficiency, not classification accuracy.